OpenStudy (inspectorjoe):

need help correcting my answers on two problems: (6,1) and (-20,35) also (60, 5) and (-20, 35)

jimthompson5910 (jim_thompson5910):

you'll need to post the full problem please

OpenStudy (inspectorjoe):

find out the distance between the pair of points. give an exact answer and where appropriate an approximation to there decimal place

jimthompson5910 (jim_thompson5910):

thanks

jimthompson5910 (jim_thompson5910):

you'll need to use the distance formula \[\Large d = \sqrt{\left(x_{2}-x_{1}\right)^2+\left(y_{2}-y_{1}\right)^2}\]

jimthompson5910 (jim_thompson5910):

for (6,1) and (-20,35) we see that x1 = 6 y1 = 1 x2 = -20 y2 = 35

OpenStudy (inspectorjoe):

yes i know but my answers are off trying to figure out why or how it should look since i can't seems to figure out why

jimthompson5910 (jim_thompson5910):

what answer did you get for the first one

OpenStudy (inspectorjoe):

yes I got that

OpenStudy (inspectorjoe):

the first one I got \[D=\sqrt{2,500}\]

OpenStudy (inspectorjoe):

the second i got \[D=\sqrt{1492}\]

jimthompson5910 (jim_thompson5910):

the first one is too big

jimthompson5910 (jim_thompson5910):

one sec

jimthompson5910 (jim_thompson5910):

For for (6,1) and (-20,35) \[\Large d = \sqrt{\left(x_{2}-x_{1}\right)^2+\left(y_{2}-y_{1}\right)^2}\] \[\Large d = \sqrt{\left(6-(-20)\right)^2+\left(1-35\right)^2}\] \[\Large d = \sqrt{\left(6+20\right)^2+\left(1-35\right)^2}\] \[\Large d = \sqrt{\left(26\right)^2+\left(-34\right)^2}\] \[\Large d = \sqrt{676+1156}\] \[\Large d = \sqrt{1832}\] Now you can simplify the radical to get... \[\Large d = \sqrt{1832}\] \[\Large d = \sqrt{4*458}\] \[\Large d = \sqrt{4}*\sqrt{458}\] \[\Large d = 2\sqrt{458}\]

jimthompson5910 (jim_thompson5910):

hopefully you can see how I got all that

OpenStudy (inspectorjoe):

the book gives us an answer of \[\sqrt{45,6.708}\]

jimthompson5910 (jim_thompson5910):

something is odd about that format

jimthompson5910 (jim_thompson5910):

are you sure it says that?

OpenStudy (inspectorjoe):

yes

OpenStudy (inspectorjoe):

question though why do you go the opposite when replacing the formula with the problem. or does it matter. meaning you did X1 - X2 and not X2-X1